A viewing screen is separate from double slit source by 1.2m the distance between the two slits is 0.03mm the second order bright fringe (m=2)is 4.5cm from the center line .(a) determine the wavelength of the light.(b) calculate the distance between adjaseet bright fringe



1
Expert's answer
2021-04-13T07:00:02-0400


Given some information



Screen is separated from double slit

D = 1.2m


Distance between two slits ( d) is 0.030 mm


Second order bight fringe is 4.5 cm from center line


"\boxed{ y_2= {2\lambda D\over d}}"


"\implies"


"\boxed{ \lambda= {y_2 d\over 2D}}"


"\boxed{ \lambda= {45\times 3 \times 10^{-3}\over 2.4}}"


"\implies"


We get the answer


(a)

"\boxed { \lambda = 562.5nm }"



Distance between adjacent bright fringe

(b)


"\boxed{\Delta y= {\lambda d\over d} = 22.5\times 10^{-3}mm}"